# Applications and resource estimates for open system simulation on a quantum computer

### v1

@@ V.3 Physical
We collect the results of our estimates on the resource estimate vs. utility plot in Fig. 6. For completeness, we also indicate where analog quantum simulators[king2021quantum] may lie. While the runtime of the universal quantum computer appears discouraging, we would like to note that our estimate is neither overly optimistic or overly pessimistic. The numbers may improve with the invention of a qubit hardware faster or more noise-resilient than the current superconducting qubit technology with $10^{-3}$ error rate and $300$ns code cycle that we used. We also expect algorithmic improvements as the open system simulation algorithms are in early stages of development. The fault-tolerant architecture itself may change as the first experimental tests of it enable its optimization beyond the conventional $T$-state distillation approach. On the other hand, the factors in big-O notation and the subleading terms, as well as the contribution of the Hamiltonian dynamics we dropped in this estimate would likely increase our numbers by several orders of magnitude. The hardware may also experience new limitations as it is scaled up to tens of millions of qubits, which are unseen in the current generation of devices. We hope our investigation pointed out the urgent open questions in the field and provided a more tempered expectation for a broader community.
@@ VI Acknowledgments
We thank Tameem Albash and Joshua Job for contributing to an early version of the draft. This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001122C0063. All material, except scientific articles or papers published in scientific journals, must, in addition to any notices or disclaimers by the Contractor, also contain the following disclaimer: Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Defense Advanced Research Projects Agency (DARPA).
Appendix A Planted solution problems
@@
A.1 Closed system: Integrable models and T-gates

### v2

@@ V.5 Physical
We collect the results of our estimates on the resource estimate vs. utility plot in Fig. 7. For completeness, we also indicate where analog quantum simulators [king2021quantum] may lie. The correct non-equilibrium behavior was observed on D-Wave, yet there are many hardware constraints preventing a true ab initio simulation. While the runtime of the universal quantum computer appears discouraging, we note that our estimate is neither overly optimistic nor overly pessimistic. The numbers may improve with the invention of a qubit hardware that is faster or more noise-resilient than the current superconducting qubit technology with $10^{-3}$ error rate and $300$ns code cycle that we used. We also expect algorithmic improvements as the open system simulation algorithms are in the early stages of development. The fault-tolerant architecture itself may change as the first experimental tests of it enable its optimization beyond the conventional T-state distillation approach. On the other hand, the factors in big-O notation and the subleading terms, as well as the contribution of the Hamiltonian dynamics we dropped in this estimate would likely increase our numbers by several orders of magnitude. The hardware may also experience new limitations as it is scaled up to tens of millions of qubits, which are unseen in the current generation of devices. We hope our investigation pointed out the urgent open questions in the field and provided a more tempered expectation for a broader community.
@@ VI Acknowledgments
We thank Tameem Albash and Joshua Job for contributing to an early version of the draft. We thank Kaitlyn Morrell, Kevin Obenland, and the rest of the team working on pyLIQTR for implementing new Hamiltonian models for us [pyLIQTR].
@@
This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001122C0063. J.M. is thankful for support from NASA Academic Mission Services, Contract No. NNA16BD14C. N.A. is a KBR employee working under Prime Contract No. 80ARC020D0010 with the NASA Ames Research Center. The United States Government retains, and by accepting the article for publication, the publisher acknowledges that the United States Government retains, a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this work or allow others to do so, for United States Government purposes. All material, except scientific articles or papers published in scientific journals, must, in addition to any notices or disclaimers by the Contractor, also contain the following disclaimer: Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Defense Advanced Research Projects Agency (DARPA).
Appendix A A sketch of an economic analysis of utility
We note that technology is definitely not enough to create a product; one also needs people and assets, as well as various other costs. While a company cannot fill the market in question without the corresponding technology, it would not be accurate to count the entire revenue as the utility of the technology. The founder who discovers the technology either invests his own money or seeks venture funding by selling shares of the company when the company is created. Only with that money can the company fill the market. Since there are various monetary figures at various stages of this process, and the revenue itself grows with time, we need a standardized method to reduce it all to a single number - utility. Note that the first stage of venture funding usually values the company relatively cheap, no matter how good the underlying technology is. We see that as the tradition of venture capital and not as the true value of the corresponding technology. We believe that the true value of technologies shows itself in acquisition events, where a big company acquires promising ideas before they become competitors. We also note that for utility estimates, the total revenue is less important than the relative increase with or without the technology. Ideally, that can be computed from the business model and an estimate of how much the technology improves the product. That being said, none of those numbers can easily be used in a unified framework, as they are rarely available. We condense all of these considerations into a factor between 0.1 and 10 that is used in our methodology. The factors given as examples in the main text are $1/2$ for nitrogen fixation alternatives and $10$ for thermoelectric devices.
A more detailed estimate of the factor $1/2$ would be as follows: first, we assume a simple model of seller optimizing profit, and the number of sales n depending linearly on price as $n=n_{0}-ep$, where e is the elasticity of demand. The result of profit optimization for production costs $c1$ and $c2$ is that the difference in revenue is $r_{2}-r_{1}=\frac{1}{2}e(c_{1}-c_{2})$. As the fertilizer market is tied with the population and the arable land, we can estimate elasticity as the remaining need for fertilizer divided by the current price since, at zero price, only the amount needed for current land and population would be purchased. The amount of fertilizer needed would remain approximately the same over a 10-year span. Finally, we estimate lower production costs $c_{2}$ from the reduction in energy consumption to fix nitrogen. Various methods have $O(1)$ difference in energy cost per unit of mass of nitrogen fixed, so we can expect a novel method to have of order 50% improvement. If it is only a 1% improvement, the estimate can be updated by directly multiplying the factors. Assuming the 50% energy savings, prices close to the production cost, and elasticity $\sim n_{0}/$current price, this estimate results in a factor $1/2$ improvement in revenue for the new nitrogen fixation technology.
In contrast, a slightly better product requiring a lot of investment into switching the manufacturing process completely would have two competing effects: rising production costs would lower the revenue, yet the improvement in product quality would increase the revenue, though the effect may be quite complex to formulate. For example, consider thermoelectric devices. While they are sold commercially today, the market is small, mostly consisting of niche cooling applications. The main application of recycling heat is mostly unrealized. It is believed that improvements in efficiency (essentially, the percentage of heat recycled) would make it viable and result in mass adoption. More specifically, the efficiency, together with other figures of merit involving lifespan, production, and maintenance cost of the thermoelectric device, determines the viability for various use cases. The demand in each use case would discontinuously jump from zero to mass adoption, but if one adds these jumps for all use cases, one could approximate the number of sales as linear in the figures of merit. Using the formalism above, that would be $n_{0}=aZT$. The change in revenue would then be $en_{0}a(ZT_{2}-ZT_{1})$. If one expects a tenfold increase in the number of sales, that would roughly translate to a tenfold increase in revenue, assuming other coefficients are $O(1)$. That illustrates that technologies with a small existing market can have much higher utility than the current or projected market size suggests. This applies only to conventionally impossible tasks, e.g., profitable heat recycling is essentially impossible with the current technology; therefore, the existing market is not a good estimate for the potential market. For most technologies, however, the task is already accomplished by the older generation of technology, and the shift in demand is not as dramatic, no matter how good the new generation of technology is.
Appendix B Planted solution problems
@@
B.1 Closed system: Integrable models and T-gates